OpenStudy (anonymous):

A scoop of ice cream, shaped like a sphere with a diameter of 6 cm, is placed in an ice cream cone with a diameter of 5cm and a height of 9 cm. Is the cone big enough to hold all the ice cream if it melts? Explain.

OpenStudy (anonymous):

The amount of ice cream when it's frozen is the same as when it is melted. so that means if we can find the volume of the sphere, we can compare it to the volume of the cone, and see if it is less than the cone's volume. Can you find the volume of the sphere of ice cream for me?

OpenStudy (anonymous):

\[V Sphere = 4/3\pi r^3\] and Volume of a Cone = \[h/3 * \pi r^2\] Remember to divide your diameter by 2 to get the radius of each of the shapes before plugging into the formulas. Compare the volumes of each to see if the cone is greater than or equal to the sphere.

OpenStudy (anonymous):

okay well can you help me through the problem

OpenStudy (anonymous):

first find the volume of the sphere by figuring out the radius of the ice cream

OpenStudy (anonymous):

I need help finding the radius

OpenStudy (anonymous):

Your diameter is 6 so if you divide by 3 your radius is 3

OpenStudy (anonymous):

okay well I need help with this whole problem can you just walk me through it?

OpenStudy (anonymous):

Plug in \[4/3 \pi 3^3\] This will give you your volume of the sphere

OpenStudy (anonymous):

Approximately 113

OpenStudy (anonymous):

Now you need to know your volume of the cone so find the radius of the base of the cone by dividing the diameter by 2.

OpenStudy (anonymous):

Knowing your height of 9 and a diameter of 5 you should plug in to the cone formula and get \[9/3 * 3.14*2.5^2\]

OpenStudy (anonymous):

This will be approximately 58.8 so, no the cone will not hold the ice cream if it all melts.